# The allosteric lever: towards a principle of specific allosteric response

### v2

@@ I.2 Full nonlinear mechanical response
$U_{2}(\mathbf{R}|\mathbf{R}^{(0)})\!=\!\frac{1}{2}\mathbf{s}^{T}\,{\textsf{S}}\,\mathbf{s}+([\mathbf{c}-\mathbf{c}^{(0)}]^{T}{\textsf{B}}-[\mathbf{s}^{(0)}]^{T}{\textsf{S}})\mathbf{s}+{\rm const},$ (6)
@@ II Response to local perturbation and allosteric lever
Figure 2: Energetics of two-step response unveils the allosteric lever. Statistics of $\delta U^{i}/\delta U^{s}$ for all source-pocket candidates during perturbation (orange) and relaxation (green) for pseudoproteins (a) and proteins (f); Energy changes $\delta U$ are normalized per-bead to allow for a comparison of networks and proteins with different sizes. The results of the true source pocket are highlighted with dashed lines at ±$\pm 1$. Alternatively we compare responses to perturbing the pocket $i$ during loading $\delta\mathbf{C}^{i}$ and relaxation $\delta\mathbf{S}^{i}$ in terms of their projections $\alpha_{k}$ onto nonzero eigenmodes of the corresponding Hessian matrix, $\mathbf{v}_{k}$ (see Eq. (12)) for pseudoproteins (b-e) and proteins (g-j) for true source pockets (bottom) and other binding-pocket candidates (top); note the logarithmic scale. To allow for a consistent comparison of structures with different sizes, the eigenmode indices $k$ are spread evenly between 1 and 10; the complete mapping onto the normalized eigenmode index is ^$\hat{\mathcal{L}}(l)$ (see Eq. (13)). Both pseudoproteins and proteins show uniform projections over the mode indices for non-allosteric pockets. In the case of true source pockets the perturbation does not couple to soft modes, whereas the response in fact predominantly couples to these, which we refer to as the allosteric lever. There is no qualitative difference between pseudoproteins and proteins.
@@
To allow for a comparison of networks and proteins with different sizes (and thus different number of eigenmodes) we divide by dimensionality $3N$ and spread the eigenmode index $l$ for each network on a logarithmic scale, yielding the normalized eigenmode index defined as
^$\hat{\mathcal{L}}(l)\equiv\log_{10}\left[\frac{9(l-6)}{3N}+1\right],\quad l=6,\ldots,3N.$ (13)
We inspect histograms of $|\alpha_{l}|$ binned as a function of ^$\hat{\mathcal{L}}(l)$ for each incremental step of the response for all pseudoproteins (Fig. 2b-e) and proteins (Fig. 2g-j). We find that in comparison to “false” pockets, the perturbation of true source pockets evidently avoids coupling to soft modes (compare Fig. 2c and Fig. 2e for pseudoproteins and Fig. 2h and Fig. 2j for proteins). Even more prominently, the concurrent relaxation upon loading true source pockets predominantly involves soft modes, whereas in the case of “false” pockets it involves all modes equally (compare Fig. 2b and Fig. 2d for pseudoproteins and Fig. 2g and Fig. 2i for proteins). This is in line with our hypothesis and confirms point 1.
@@ III.1 Avoided eigenvalue crossings and origin of non-linear response
We observe a strong correlation between the degree of non-linearity and the number of avoided crossings $N_{\rm cxs}$ relative to the rank of the matrix H (see Fig. 4c and 4d for pseudoproteins and proteins, respectively; the two highly non-linear “outliers” in Fig. 4c are discussed in Appendix G). This correlation is not surprising; avoided crossings are known to occur when vibrations (in classical as well as quantum systems [101, 102, 103]) become strongly coupled by an external perturbation.
@@
Figure 4: Nonlinear response and avoided crossings. (a-b) Eigenvalues $\omega_{l}(k)$ of the instantaneous Hessian at the $k$-th step of the response of (a) a self-trained pseudoprotein and (b) HSA protein (PDB ID: 2bxg [68]); red crosses indicate the estimated positions of avoided crossings and the 3D insets show representations of the respective network. (c-d) Scatter plots depicting a clear correlation between the number of avoided crossings (normalized by the total number of internal eigenvectors), $N_{\rm cxs}/{\rm rank({\textsf{H}})}$, and the nonlinearity of the response quantified as the rotation of the eigenvectors of the instantaneous Hessian during the response with respect to the initial (resting) configuration averaged over the response trajectory, $\langle\beta\rangle$ in Eq. (16), for pseudoproteins (c) and proteins (d). The two highly nonlinear outliers in (d) (both structures of the same motor protein (Myosin V), performing larger motions) are discussed in Appendix G.
@@ IV Specificity of allosteric responses
such that the desired/optimal response has relative magnitude and specificity one, $\Delta_{*}=\mu_{*}=1$. Triangle inequalities between the norms in Eqs. (17-18) (see  [^3] We have, setting $\mathbf{a}\equiv\mathbf{T}_{i}^{\rm fin}-\mathbf{T}^{0}$, $\mathbf{b}\equiv\mathbf{T}_{i}^{\rm fin}-\mathbf{T}_{*}^{\rm fin}$ and $\mathbf{c}\equiv\mathbf{T}_{*}^{\rm fin}-\mathbf{T}^{0}$, that $\mathbf{a}=\mathbf{b}+\mathbf{c}$ and hence $|\mathbf{a}|\leq|\mathbf{b}|+|\mathbf{c}|$, $|\mathbf{c}|\geq||\mathbf{b}|-|\mathbf{a}||$ and $|\mathbf{b}|\geq||\mathbf{c}|-|\mathbf{a}||$; the squeeze bound Eq. (20) follows. Note also that rotations and translations preserve metric properties [155]. for details) confine all responses to the domain
@@
Figure 5: Optimal source pockets efficiently transmit energy and yield large and specific allosteric responses. Joint frequency histograms of magnitude $\Delta$ and specificity $\mu$, $p(\Delta,\mu)$ (defined in Eqs. (17) and (19), respectively) and their marginals, $p(\Delta)$ and $p(\mu)$, for adjacent pairs of beads on the surface of pseudoproteins (a) and proteins (b). The histograms are evaluated for $\Delta_{i}$ and $\mu_{i}$ for all pairs $i$ of surface beads normalized with respect to the given network, and the statistics is determined over all networks. The black lines denote the bounds imposed by the three triangle inequalities between the norms in Eqs. (17-18). Green arrows point to the ideal response at $(1,1)$. The latter may not necessarily be reached in practical applications (e.g. in drug design); in this case source candidates nearest to the optimum are to be considered. Inset: magnification of the desired region. (g-j) Probability densities of energy changes $\delta U^{i}$ upon perturbing site $i$ relative to that of perturbing the true source $\delta U^{s}$ during loading (orange shaded) and relaxation (green shaded) substeps conditioned on a large response $\Delta$ (panels (g) and (i)) and a highly specific response $\mu$ (panels (h) and (j)); green and orange lines correspond to unconditional probability densities from Fig. 2 and unveil a clear shift towards the tails. The response was divided in 100 increments for both, pseudoproteins and proteins.
@@ VI Evolutionary conservation of allosteric transmission channels
We find that true source-pocket residues are conserved more than the average, with a few exceptions (see Fig. 7a). The beads we predict as potential source beads are conserved with the same, or an even slightly higher rate (see Fig. 7b). This may be interpreted as that not all of the residues in the source pocket are important for allostery even though the ligand binds to them. This idea is strengthened in context of findings in Appendix A.1, where we find that not all of the source beads directly contribute to the energy input. When constraining combinations of smaller subsets (i.e. randomly chosen triplets) of beads within the source pocket, we instead find that some beads do not lead to an energy change during the constraining step but instead only during the relaxation step, i.e. when the truly critical beads follow their displacement. Moreover, sites that are involved in efficient energy transmission—beads connected by a spring that contributes strongly to the energy change during relaxation—are as well conserved more than the average (see Fig. 7c). This confirms point 4 of our allosteric-lever hypothesis. Notably, the identified efficient-energy-transmission beads are potentially related to the so-called “dynamic allosteric residue coupling (DARC) spots”, whose mutations were hypothesized to cause changes in the allosteric networking [115].
@@ Appendix D Removing rigid-body motions
Consider the square matrix ${\textsf{M}}\in\mathbbm{R}^{m\times m}$ which does not have full rank $r$, its nullity being $n$, such that $n+r=m$. The nullspace is $\mathcal{N}({\textsf{M}})$ with dimension $n$ and its row space (or range) is $\mathcal{R}({\textsf{M}}^{T})$ with dimension $r$. All vectors in $\mathcal{N}({\textsf{M}})$ are orthogonal to all vectors in $\mathcal{R}({\textsf{M}}^{T})$ Therefore, if we add a matrix given by an outer product of a basis vector $\mathbf{x}$ from the nullspace to M, i.e. $\tilde{{\textsf{M}}}\equiv{\textsf{M}}+\mathbf{x}\otimes\mathbf{x}^{T}$, we increase the rank of M by one. Applying this for each basis vector of $\mathcal{N}({\textsf{M}})$ [^4] Standard eigendecomposition routines are not ideal for a fast calculation of the eigenvectors that correspond to the 6 zero eigenvalues of H. Using a shift and invert spectral transformation from ARPACK [156], where we choose the number of eigenvalues we are interested in and provide a starting point for the search, is considerably faster. The corresponding implementation in scipy [157] does not always converge for eigenvalues that are of the same magnitude as the numerical precision (i.e. zero eigenvalues). We overcome this issue by adding a small noise to the matrix whose nullspace is to be determined. gives M full rank [154]. This, however, does not change any result where the matrix is applied to a vector $\mathbf{y}$ from the row space $\mathcal{R}({\textsf{M}}^{T})$, since

### v3

@@ I.2 Full nonlinear mechanical response
$U_{2}(\mathbf{R}|\mathbf{R}^{(0)})\!=\!\frac{1}{2}\mathbf{s}^{T}\,{\textsf{S}}\,\mathbf{s}+([\mathbf{c}-\mathbf{c}^{(0)}]^{T}{\textsf{B}}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T}}-[\mathbf{s}^{(0)}]^{T}{\textsf{S}})\mathbf{s}+{\rm const},$ (6)
@@ II Response to local perturbation and allosteric lever
Figure 2: Energetics of two-step response unveils the allosteric lever. Statistics of $\delta U^{i}/\delta U^{s}$ for all source-pocket candidates during perturbation (orange) and relaxation (green) for pseudoproteins (a) and proteins (f); Energy changes $\delta U$ are normalized per-bead to allow for a comparison of networks and proteins with different sizes. The results of the true source pocket are highlighted with dashed lines at ±$\pm 1$. Alternatively we compare responses to perturbing the pocket $i$ during loading $\delta\mathbf{C}^{i}$ and relaxation $\delta\mathbf{S}^{i}$ in terms of their projections $\alpha_{k}$ onto nonzero eigenmodes of the corresponding Hessian matrix, $\mathbf{v}_{k}$ (see Eq. (12)) for pseudoproteins (b-e) and proteins (g-j) for true source pockets (bottom) and other binding-pocket candidates (top); note the logarithmic scale. To allow for a consistent comparison of structures with different sizes, the eigenmode indices $k$ are spread evenly between 1 and 10 and shown on the double-logarithmic scale for better visibility; the complete mapping onto the normalized eigenmode index is ^$\hat{\mathcal{L}}(l)$ (see Eq. (13)). Both pseudoproteins and proteins show uniform projections over the mode indices for non-allosteric pockets. In the case of true source pockets the perturbation does not couple to soft modes, whereas the response in fact predominantly couples to these, which we refer to as the allosteric lever. There is no qualitative difference between pseudoproteins and proteins.
@@
To allow for a comparison of networks and proteins with different sizes (and thus different number of eigenmodes) we divide by dimensionality $3N$ and map the eigenmode index $l$ for each network uniformly onto the interval $[1,10]$, yielding the normalized eigenmode index defined as
^${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\mathcal{L}}(l)\equiv\left[9\frac{(l-6)}{(3N-6)}+1\right],\quad l=6,\ldots,3N.}$ (13)
By construction ^$\hat{\mathcal{L}}(l)$ lies between $1$ and $10$ for all structures. We inspect histograms of $|\alpha_{l}|$ binned as a function of ^$\hat{\mathcal{L}}(l)$ for each incremental step of the response for all pseudoproteins (Fig. 2b-e) and proteins (Fig. 2g-j). We find that in comparison to “false” pockets, the perturbation of true source pockets evidently avoids coupling to soft modes (compare Fig. 2c and Fig. 2e for pseudoproteins and Fig. 2h and Fig. 2j for proteins). Even more prominently, the concurrent relaxation upon loading true source pockets predominantly involves soft modes, whereas in the case of “false” pockets it involves all modes equally (compare Fig. 2b and Fig. 2d for pseudoproteins and Fig. 2g and Fig. 2i for proteins). This is in line with our hypothesis and confirms point 1.
@@ III.1 Avoided eigenvalue crossings and origin of non-linear response
We observe a strong (not necessarily linear) correlation between the degree of non-linearity and the number of avoided crossings $N_{\rm cxs}$ relative to the rank of the matrix H (see Fig. 4c and 4d for pseudoproteins and proteins, respectively; the two highly non-linear “outliers” in Fig. 4c are discussed in Appendix G). We quantify this (generally nonlinear) correlation by means of the (sample) distance correlation, ${\rm dCor}$ [101]. That is, we consider the tuples $V_{1}\equiv\{\frac{N_{\text{cxs}}}{\text{rank}(\mathbf{H})}(i)\}_{i=1,n_{s}}$ and $V_{2}\equiv\{1-\langle\beta\rangle(i)\}_{i=1,n_{s}}$ where $n_{s}$ denotes, respectively, the number of protein and pseudoprotein structures considered in the analysis, and define pairwise distances $D^{\alpha}_{kl}\equiv|V_{\alpha}(k)-V_{\alpha}(l)|$ for $\alpha=1,2$ in the Euclidean norm ||$\cdot$|| ⋅ |. We further define the (doubly centered) symmetric distance matrices ${\textsf{M}}_{\alpha}$ with elements
$({\textsf{M}}_{\alpha})_{ij}=D^{\alpha}_{ij}-\frac{1}{n_{s}}\sum_{k=1}^{n_{s}}(D^{\alpha}_{ik}+D^{\alpha}_{kj})+\frac{1}{n_{s}^{2}}\sum_{k,l=1}^{n_{s}}D^{\alpha}_{kl},$ (17)
as well as (scalar-valued) distance covariances
${\rm dCov}(V_{1},V_{2})\equiv\frac{1}{n_{s}^{2}}{\rm Tr}\,\left({\textsf{M}}_{1}{\textsf{M}}_{2}^{T}\right)$ (18)
and distance variances $\text{dVar}(V_{\alpha})={\rm dCov}(V_{\alpha},V_{\alpha})$ that enter the definition of the distance correlation
${\rm dCor}(V_{1},V_{2})\equiv\frac{{\rm dCov}(V_{1},V_{2})}{\sqrt{\text{dVar}(V_{1})\text{dVar}(V_{2})}}.$ (19)
We have $0\leq{\rm dCor}(V_{1},V_{2})\leq 1$ and ${\rm dCor}(V_{1},V_{2})=0$ if and only if $V_{1}$ and $V_{2}$ are independent [101], so ${\rm dCor}(V_{1},V_{2})>0$ indeed quantifies statistical dependence between $V_{1}$ and $V_{2}$. We find ${\rm dCor}\simeq 0.62$ for the proteins and ${\rm dCor}\simeq 0.83$ for pseudoproteins, respectively, quantitatively confirming substantial correlations between the degree of non-linearity and the number of avoided crossings $N_{\rm cxs}$.
The observed correlation is not surprising; avoided crossings are known to occur when vibrations (in classical as well as quantum systems [102, 103, 104]) become strongly coupled by an external perturbation.
@@
Figure 4: Nonlinear response and avoided crossings. (a-b) Eigenvalues $\omega_{l}(k)$ of the instantaneous Hessian at the $k$-th step of the response of (a) a self-trained pseudoprotein and (b) HSA protein (PDB ID: 2bxg [68]); red crosses indicate the estimated positions of avoided crossings and the 3D insets show representations of the respective network. (c-d) Scatter plots depicting a clear correlation between the number of avoided crossings (normalized by the total number of internal eigenvectors), $N_{\rm cxs}/{\rm rank({\textsf{H}})}$, and the nonlinearity of the response quantified as the rotation of the eigenvectors of the instantaneous Hessian during the response with respect to the initial (resting) configuration averaged over the response trajectory, $\langle\beta\rangle$ in Eq. (16), for pseudoproteins (c) and proteins (d). The two highly nonlinear outliers in (d) (both structures of the same motor protein (Myosin V), performing larger motions) are discussed in Appendix GThe magnitude of the nonlinear correlation is quantified in terms of the distance correlation ${\rm dCor}$ defined in Eq. (19).
@@ IV Specificity of allosteric responses
such that the desired/optimal response has relative magnitude and specificity one, $\Delta_{*}=\mu_{*}=1$. Triangle inequalities between the norms in Eqs. (20-21) (see  [^3] We have, setting $\mathbf{a}\equiv\mathbf{T}_{i}^{\rm fin}-\mathbf{T}^{0}$, $\mathbf{b}\equiv\mathbf{T}_{i}^{\rm fin}-\mathbf{T}_{*}^{\rm fin}$ and $\mathbf{c}\equiv\mathbf{T}_{*}^{\rm fin}-\mathbf{T}^{0}$, that $\mathbf{a}=\mathbf{b}+\mathbf{c}$ and hence $|\mathbf{a}|\leq|\mathbf{b}|+|\mathbf{c}|$, $|\mathbf{c}|\geq||\mathbf{b}|-|\mathbf{a}||$ and $|\mathbf{b}|\geq||\mathbf{c}|-|\mathbf{a}||$; the squeeze bound Eq. (23) follows. Note also that rotations and translations preserve metric properties [Fogolari]. for details) confine all responses to the domain
@@
Figure 5: Optimal source pockets efficiently transmit energy and yield large and specific allosteric responses. Joint frequency histograms of magnitude $\Delta$ and specificity $\mu$, $p(\Delta,\mu)$ (defined in Eqs. (20) and (22), respectively) and their marginals, $p(\Delta)$ and $p(\mu)$, for adjacent pairs of beads on the surface of pseudoproteins (a) and proteins (b). The histograms are evaluated for $\Delta_{i}$ and $\mu_{i}$ for all pairs $i$ of surface beads normalized with respect to the given network, and the statistics is determined over all networks. The black lines denote the bounds Eq. (23) imposed by the three triangle inequalities between the norms in Eqs. (20-21). Green arrows point to the ideal response at $(1,1)$. The latter may not necessarily be reached in practical applications (e.g. in drug design); in this case source candidates nearest to the optimum are to be considered. Inset: magnification of the desired region. (g-j) Probability densities of energy changes $\delta U^{i}$ upon perturbing site $i$ relative to that of perturbing the true source $\delta U^{s}$ during loading (orange shaded) and relaxation (green shaded) substeps conditioned on a large response $\Delta$ (panels (g) and (i)) and a highly specific response $\mu$ (panels (h) and (j)); green and orange lines correspond to unconditional probability densities from Fig. 2 and unveil a clear shift towards the tails. The response was divided in 100 increments for both, pseudoproteins and proteins.
@@ VI Evolutionary conservation of allosteric transmission channels
We find that true source-pocket residues are conserved more than the average, with a few exceptions (see Fig. 7a). The beads we predict as potential source beads are conserved with the same, or an even slightly higher rate (see Fig. 7b). This may be interpreted as that not all of the residues in the source pocket are important for allostery even though the ligand binds to them. This idea is strengthened in context of findings in Appendix A.1, where we find that not all of the source beads directly contribute to the energy input. When constraining combinations of smaller subsets (i.e. randomly chosen triplets) of beads within the source pocket, we instead find that some beads do not lead to an energy change during the constraining step but instead only during the relaxation step, i.e. when the truly critical beads follow their displacement. Moreover, sites that are involved in efficient energy transmission—beads connected by a spring that contributes strongly to the energy change during relaxation—are as well conserved more than the average (see Fig. 7c). This confirms point 4 of our allosteric-lever hypothesis. Notably, the identified efficient-energy-transmission beads are potentially related to the so-called “dynamic allosteric residue coupling (DARC) spots”, whose mutations were hypothesized to cause changes in the allosteric networking [116]. This feature may also be a coincidence, as mutations in proteins are selected for many reasons and ENM do not explicitly encode any chemical specificity.
@@ Appendix D Removing rigid-body motions
Consider the square matrix ${\textsf{M}}\in\mathbbm{R}^{m\times m}$ which does not have full rank $r$, its nullity being $n$, such that $n+r=m$. The nullspace is $\mathcal{N}({\textsf{M}})$ with dimension $n$ and its row space (or range) is $\mathcal{R}({\textsf{M}}^{T})$ with dimension $r$. All vectors in $\mathcal{N}({\textsf{M}})$ are orthogonal to all vectors in $\mathcal{R}({\textsf{M}}^{T})$ Therefore, if we add a matrix given by an outer product of a basis vector $\mathbf{x}$ from the nullspace to M, i.e. $\tilde{{\textsf{M}}}\equiv{\textsf{M}}+\mathbf{x}\otimes\mathbf{x}^{T}$, we increase the rank of M by one. Applying this for each basis vector of $\mathcal{N}({\textsf{M}})$ [^4] Standard eigendecomposition routines are not ideal for a fast calculation of the eigenvectors that correspond to the 6 zero eigenvalues of H. Using a shift and invert spectral transformation from ARPACK [Lehoucq1998ARPACKUG], where we choose the number of eigenvalues we are interested in and provide a starting point for the search, is considerably faster. The corresponding implementation in scipy [scipy] does not always converge for eigenvalues that are of the same magnitude as the numerical precision (i.e. zero eigenvalues). We overcome this issue by adding a small noise to the matrix whose nullspace is to be determined. gives M full rank [155]. This, however, does not change any result where the matrix is applied to a vector $\mathbf{y}$ from the row space $\mathcal{R}({\textsf{M}}^{T})$, since